# Time Evolution in Canonical Quantum Gravity is Trivial

### v1

@@ Abstract
The Wheeler-DeWitt (WdW) equation does not describe any explicit time evolution of the wave function, and somehow related to this issue, there is no natural way of defining an invariant inner product that provides a viable probability interpretation. We show that both of these difficulties are solved in a covariant canonical formulation of general relativity where the configuration space is extended by introducing the embedding coordinates as dynamical variables. The formalism describes the evolution of the wave function from one spacelike slice to another, but as in the case of spatial diffeomorphisms this is simply implemented by a coordinate change in the wave function. We demonstrate how the time equation disappears after gauge fixing that removes the embedding coordinates. These findings indicate that the time evolution is trivial in a background independent formulation of quantum gravity.
@@
There are various proposals to solve the problem of time in quantum gravity, however a satisfactory solution has not yet be found (see k1 for an extensive review). For example, a standard approach is to deparametrize the theory by gauge fixing, after which a field (or a combination of the fields) in the theory becomes an emergent time parameter that keeps track of time evolution (the remaining degrees of freedom are quantized in the standard way). Although such a procedure seems straightforwardly applicable, e.g., in the path integral quantization bar1 ; bar2 , the gauge fixing generally breaks down and consequently the field ceases to be an honest time variable. Only in some special models one can find a suitable variable that can play the role of time. For example, the scale factor of the universe can be used as an emergent time in a flat minisuperspace quantum cosmology of a self-interacting scalar field having a positive definite potential; yet it fails to be an honest time variable if the potential is not positive definite ak1 ; ak2 . Thus, deparametrization does not offer a satisfactory general solution to the problem of time and other proposals have similar issues.
@@
Figure 1: An initial spacelike constant-$t$ slice and its deformation along the normal vector field in a local region R. The deformation should yield a new speclike slice.
@@
where GF and FP stands for the gauge fixing and Faddeev-Popov terms. One can show that (10) does not change under the deformation $\delta X^{\mu}=\epsilon(y^{i})n^{\mu}$ hence the inner product is invariant under time translations, see aks2 for details. The real physical degrees of freedom can be identified as the metric $h_{ij}$ modulo coordinate transformations and obviously $|\Psi[X^{\mu},h_{ij}]|^{2}$ provides a meaningful probability distribution.
@@
where $\det(\gamma_{ij})=1$. The condition $\Phi\Psi=0$ gives a functional differential equation which can be used to fix the scale factor $a$ dependence of $\Psi$. In general, there are two independent solutions (for each $a(y^{i})$ variable) since $\Phi\Psi=0$ is second order in functional derivatives but only one solution is expected to be normalizable under (10) as $a(y^{i})\to\infty$ (as discussed in aks2 , the nomalizability may also require extra conditions in the theory, e.g. there must exist a negative cosmological constant in the simplest minisuperspace model.)
@@
We note that the structure of the theory completely changes when one expands around a classical background solution. In ak3 , we have systematically studied how such an expansion in the action can be carried out for the perturbations around the background and showed that there appears an honest/non-vanishing Hamiltonian and additional first class constraints. Structure-wise, the theory becomes similar to a typical Yang-Mills gauge theory (that have infinitely many terms) and the problem of time is bypassed because the background solution provides a reference for time evolution. In the context of the WdW equation, one can imagine a wavefunction of the form $\Psi=\Psi_{cl}\,\psi$, where $\Psi_{cl}$ is localized around the given classical background. Then, it is possible to show under certain approximations that the time independence of the full wavefunction $\partial\Psi/\partial t=0$ together with the WdW equation yield a Schrödinger equation for $\psi$ ak4 . As a result, there emerges a significant difference between the background independent vs. dependent approaches to quantum gravity in terms of how time is dealt with.
Finally, in the absence of dynamical time evolution, it is not meaningful to define some of the observables such as scattering amplitudes, however one can still apply the conventional rules of quantum mechanics, for example, in calculating the expectation values of field operators in a specified state. The situation is similar to a totally isolated system which is prepared in an energy eigenstate, where the Schrödinger equation becomes trivial. Maybe not surprisingly, the canonical quantization implies that the universe should be considered as an isolated system having zero total energy.

### v4

@@ Abstract
The Wheeler-DeWitt (WdW) equation does not describe any explicit time evolution of the wave function, and somehow related to this issue, there is no natural way of defining an invariant inner product that provides a viable probability interpretation. We show that both of these difficulties are solved in a covariant canonical formulation of general relativity where the configuration space is extended by introducing the embedding coordinates as dynamical variables. The formalism describes the evolution of the wave function from one spacelike slice to another, but as in the case of spatial diffeomorphisms this is simply implemented by a coordinate change in the wave function. We demonstrate how the time evolution equation disappears after gauge fixing that removes the embedding coordinates. These findings indicate that the time evolution is trivial in a background independent formulation of quantum gravity.
@@
There are various proposals to solve the problem of time in quantum gravity, however a satisfactory solution has yet to be found (see k1 for an extensive review). For example, a standard approach is to deparametrize the theory by gauge fixing, after which a field (or a combination of the fields) in the theory becomes an emergent time parameter that keeps track of time evolution (the remaining degrees of freedom are quantized in the standard way). Although such a procedure seems straightforwardly applicable, e.g., in the path integral quantization bar1 ; bar2 , the gauge fixing generally breaks down and consequently the field ceases to be an honest time variable. Only in some special models one can find a suitable variable that can play the role of time. For example, the scale factor of the universe can be used as an emergent time in a flat minisuperspace quantum cosmology of a self-interacting scalar field having a positive definite potential; yet it fails to be an honest time variable if the potential is not positive definite ak1 ; ak2 . Thus, deparametrization does not offer a satisfactory general solution to the problem of time and other proposals have similar issues.
@@
The second and the third constraints in (8) generate the symmetry (4). It is possible to rewrite them (after some algebra) in the following form
$P_{\mu}+2P^{ij}\,D_{i}\left(\partial_{j}X^{\nu}\,g_{\mu\nu}\right)=0,$ (9)
which can be shown to generate (5). The former can be obtained from (9) by contracting with $\partial_{i}X^{\mu}$ and $n^{\mu}$. Although their generators are related by a (field dependent) contraction, it is crucial to note that (4) and (5) are two different symmetries; the former is generated by a world-volume vector $k^{\alpha}$ as its parameter and the latter is generated by a spacetime vector $l^{\mu}$. Note also that in general a symmetry generator simply multiplied by a field dependent term will not be a symmetry generator anymore; this happens in our case since there are actually two different symmetries.
@@
Figure 1: An initial spacelike constant-$t$ slice and its deformation along the normal vector field in a local region R. The deformation should yield a new spacelike slice.
@@
It is important to emphasize that the operator $n^{\mu}P_{\mu}$ generates a real deformation in the spacetime like in Fig. 1 and does not simply correspond to a repartmentalization of the same initial surface since $P_{\mu}$ generates the translation of the spacetime coordinate $X^{\mu}$. A possible factor which may mislead one to think on the contrary is $\delta\gamma_{\alpha\beta}=0$ under (5), however this simply means the brane in our brane-anology is not dynamical. Indeed, as shown in aks1 , (10) yields the standard Schrödinger equation for a self-interacting scalar field where $n^{\mu}P_{\mu}$ becomes the operator $\partial/\partial t$. As pointed out above, the main difference of general relativity is the extra constraint $\Phi=0$, yet the kinematics are completely identical in the two theories.
@@
where GF and FP stands for the gauge fixing and Faddeev-Popov terms (as in the standard Yang-Mills theory, this procedure should make (11) finite by removing the infinities arising from the gauge symmetry (iv)). One can show that (11) does not change under the deformation $\delta X^{\mu}=\epsilon(y^{i})n^{\mu}$ hence the inner product is invariant under time translations, see aks2 for details. The real physical degrees of freedom can be identified as the metric $h_{ij}$ modulo coordinate transformations and obviously $|\Psi[X^{\mu},h_{ij}]|^{2}$ provides a meaningful probability distribution.
@@
where $\det(\gamma_{ij})=1$ (this decomposition is irreducible under spatial diffeomorphisms, i.e. $a$ and $\gamma_{ij}$ do not mix with each other). The condition $\Phi\Psi=0$ gives a functional differential equation of the form
$\frac{\delta^{2}}{\delta a(y^{i})^{2}}\Psi=...$ (13)
which can be used to fix the scale factor $a$ dependence of $\Psi$. In general, there are two independent solutions (for each $a(y^{i})$ variable) since $\Phi\Psi=0$ is second order in functional derivatives but only one solution is expected to be normalizable under (11) as $a(y^{i})\to\infty$. As discussed in aks2 , the nomalizability may also require extra conditions in the theory, e.g. there must exist a negative cosmological constant in the simplest minisuperspace model. Also, it is crucial that the scale factor ranges in the half-line $0<a<\infty$ otherwise for a full range (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) it would be impossible to normalize the wavefunction (one should also be careful about the regularity as $a\to 0$ but it looks like the path integral measure in (11) yields enough positive powers of $a$ so that no problem arises in that limit aks2 ).
@@
We note that the structure of the theory completely changes when one expands around a classical background solution. In ak3 , we have systematically studied how such an expansion in the action can be carried out for the perturbations around the background and showed that there appears an honest/non-vanishing Hamiltonian and additional first class constraints. Structure-wise, the theory becomes similar to a typical Yang-Mills gauge theory (that have infinitely many terms) and the problem of time is bypassed because the background solution provides a reference for time evolution. In the context of the WdW equation, one can imagine a wavefunction of the form $\Psi=\Psi_{cl}\,\psi$, where $\Psi_{cl}$ is localized around the given classical background. Then, it is possible to show under certain approximations that the time independence of the full wavefunction $\partial\Psi/\partial t=0$ together with the WdW equation yield a Schrödinger equation for $\psi$ (it is also possible to utilize a WKB approximation after which a time evolution emerges with respect to the embedding coordinates viewed as a Gaussian reference fluid, see mm ; mma ). As a result, there emerges a significant difference between the background independent vs. dependent approaches to quantum gravity in terms of how time is dealt with.
The quantization procedure proposed in this paper is by no means unique and there are obviously different approaches which may not yield equivalent results due to the nonlinear structure of general relativity. As pointed out above, in certain cases quantizing the true physical degrees of freedom after gauge fixing (deparametrization) is another viable alternative, which yields a time variable and an evolving wave-function, see e.g. ak1 ; ek1 for examples. It is not difficult to see that the two quantizations, deparametrization and WdW, are inequivalent. For example, in a simple scalar field minisuperspace model which has two variables $a$ and $\phi$, the deparametrized theory yields a wavefunction $\psi(a,\phi)$, where $a$ plays the role time (hence it is non-dynamical) and $\psi(a,\phi)$ gives a probability distribution for $\phi$ ak1 . On the other hand, in the WdW quantization as we discuss here, the wave-function $\Psi(a,\phi)$ yields a probability distribution over the dynamical variables $a$ and $\phi$, and there is no time evolution. In the pure minisuperspace model without any matter, the deparametrized theory is pure gauge woodard but the WdW quantization yields a single state aks2 , again showing the inequivalence of the two methods. It would not be surprising to see that in other approaches like loop quantum gravity ek2 or refined algebraic quantization ek3 , one reaches different conclusions.
Finally, in the absence of dynamical time evolution, it is not meaningful to define some of the observables such as scattering amplitudes, however one can still apply the conventional rules of quantum mechanics, for example, in calculating the expectation values of field operators in a specified state. The situation is similar to a totally isolated system which is prepared in an energy eigenstate where the Schrödinger equation becomes trivial. Maybe not surprisingly, the canonical quantization implies that the universe should be considered as an isolated system having zero total energy. It would be interesting to develop the formalism in the presence of boundary terms, particularly for asymptotically flat spaces with finite ADM masses. In such cases, one would expect to obtain a nonzero Hamiltonian for the asymptotic fields, leading to a nontrivial time evolution of the states involving asymptotic variables. We hope to address this problem in a future publication.